OpenStudy (anonymous):

Find Solutions in interval "[0,2pi)"; cos^2 x + 2cosx + 1 = 0

OpenStudy (anonymous):

I don't so much need the answer, just how to get it. I don't fully understand the concept of intervals and how to find it by applying this "[0,2pi)".

OpenStudy (anonymous):

Trig functions are periodic, meaning that they will go on forever repeating the same kinds of values over and over. When they give you an interval, they are saying "just look in this neighborhood". When they give you 0 to 2pi, it is kind of nice because that is where you are used to looking.

OpenStudy (anonymous):

Where those lines are crossing the x-axis on his graph...would those count as intervals?

OpenStudy (anonymous):

They can, but the interval defined by your problem says to look from 0 (the y-axis) over to 2pi (the end of one cycle).

OpenStudy (anonymous):

Ok, thank you. =) I have a poor understanding of the unit circle in general so it's nice to learn SOMETHING about it

OpenStudy (anonymous):

They limited your interval (from being infinity) , because that would have given an infinite number of answers. They just wanted the standard answer that you would get if you looked "in the neigborhood" of where the function started.

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

I hate to bother you further about the same concept, but I've never seen an equation of this caliber. 7tan^3 x - 21tan x = 0. Same coordinants; what can you get from that?

OpenStudy (anonymous):

I would start by factoring out 7 tan x. Then set each factor to zero.

OpenStudy (anonymous):

Alrighty

OpenStudy (anonymous):

._. I'm trying to factor it out replacing tan with "a" but I'm having a bit of trouble. Do you have any hints or tips for me?

OpenStudy (anonymous):

my first step looked like this\[7\tan x(\tan ^{2}x-3)=0\]

OpenStudy (anonymous):

These kinds of problems are where algebra and trig collide. You look at whole trig functions as variables in a sense. Once we factor out... set each factor to 0.\[7\tan x=0\]and\[\tan ^{2}x-3=0\]

OpenStudy (anonymous):

I only use a substitution if my factoring gets really hairy... like if it takes the form \[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

How would I solve for x with "tan" in the equation?

OpenStudy (anonymous):

inverse tangent... I had a teacher force me to learn my unit circle including tangents... so I think of these as "nice" answers. \[\tan ^{2}x=3\]\[\tan x=\pm \sqrt{3}\]\[\tan^{-1} (\tan x)=\tan^{-1} (\pm \sqrt{3})\]\[x=\pm \frac{\pi}{3}\]

OpenStudy (anonymous):

cos^2 x + 2cosx +1=(cosz+1)^2=0 then cosx=-1 ans=[pi,3pi,5pi,...,(2n+1)pi give a medal plz

OpenStudy (anonymous):

lol force? And I know 4pi/3 has to be an interval...would just "pi" we one as well? How can we test this? (I have a lot of questions about this, it's ok if you leave)

OpenStudy (anonymous):

Ah... -pi/3 wasn't in the interval was it.... from 0 to 2pi.

OpenStudy (anonymous):

You've got it... 4pi/3 is the other one... perfect! And I think that pi would be one as well as zero. You are getting the hang of this interval thing!

OpenStudy (anonymous):

You could plug these numbers into your original equation to verify... just plug them in for x and see if you get zero.

OpenStudy (anonymous):

=p Is that all the possible intervals? I can't think of anymore

OpenStudy (anonymous):

My pc is doing strange things... forgive me if I lag. Remember that interval is not the answer it is simply the place where you are looking for answers.

OpenStudy (anonymous):

Given the interval from zero (inclusive) to 2pi (non-inclusive), your answers should be 0, pi/3, pi, and 4pi/3... the next one would be 2pi, but that would be outside of the interval asked about.

OpenStudy (anonymous):

ooh I think there are some hiding... remember when I said -pi/3... even though it is not in your given interval... it will produce solutions at 2pi/3 and 7pi/3.... I need to get my calculator. :)

OpenStudy (anonymous):

Hmm...im not sure about 2pi/3, but wouldnt 7pi/3 be outside our [0,2pi) boundaries?

OpenStudy (anonymous):

you are correct... brain fart... I meant 5pi/3

OpenStudy (anonymous):

ok... so here is a way to gain a little insight.. do you have a graphing calculator?

OpenStudy (anonymous):

No, but I can try to find an online one =o

OpenStudy (anonymous):

I will give my final answer now :p {0, pi/3, 2pi/3, pi, 4pi/3, 5pi/3} What do you know... every pi/3.

OpenStudy (anonymous):

Graph your original equation without the equals zero. Every x-axis crossing between [0, 2pi) is a solution.

OpenStudy (anonymous):

And that's all within [0,2pi)? The > 2 digits are throwing me off

OpenStudy (anonymous):

my answer is true in your domain just pi can be a answer

OpenStudy (anonymous):

fractions can make things look a little strange... if we are looking at pi/3's then 6pi/3 is 2pi and 5pi/3 is therefore the largest pi/3 we could use.

OpenStudy (anonymous):

mahdi... we are on a different problem from the original posting

OpenStudy (anonymous):

This one was a little rough... every time I thought I had it, there was more to it. Looking back... I would say that our original three answers (without considering interval) were -pi/3, 0, and pi/3. Due to the fact that a tangent has a period of pi, these solutions repeat every pi on the interval. If there was not a given interval... the solution would occur every pi/3 or generally x=npi/3 where n is an integer. The values of this that fall within our interval of [0,2pi) are {0, pi/3, 2pi/3, pi, 4pi/3, an 5pi/3}. whew!

OpenStudy (anonymous):

If you view a graph, you will notice that the graph crosses at every one of these locations.

OpenStudy (anonymous):

xp Goodness. So what did you put into your calculator? I wont ask you anything else, I know this is a lot

OpenStudy (anonymous):

\[7\tan ^{3}x-21\tan x\]

OpenStudy (anonymous):

Again, thank you so much for your help =) I think I'll post one more question (this one has a -7 where our 0 was, oh gawd) and call it a night

OpenStudy (anonymous):

You're very welcome. I like it when people are curious and really seek to understand more than just their one answer.

OpenStudy (anonymous):

^_^ that would explain you sticking around so long even after I already gave you a medal